OpenStudy (anonymous):

Can someone help me with interest that is compounded continuously and monthly?

OpenStudy (alexwee123):

yes what about it?

OpenStudy (anonymous):

Which formula would you use for each? I know that one is A=P(1+r/t)^n*t, and the other is A=Pe^rt, I think. But I am not sure.

OpenStudy (alexwee123):

A=Pe^rt is compounded continiuosly

OpenStudy (alexwee123):

compounded monthly= A=P(1+r/t)^n*t

OpenStudy (alexwee123):

just remember that n=12 since there are 12 months in a year

OpenStudy (alexwee123):

and it would be A=P(1+r/n)^n*t not A=P(1+r/t)^n*t

OpenStudy (alexwee123):

sorry i didn't see that

OpenStudy (anonymous):

Okay, thanks! Does e=2.718, or is that found in the equation?

OpenStudy (alexwee123):

e=2.718

OpenStudy (anonymous):

Thank you so much! I really appreciate it.

OpenStudy (alexwee123):

np:)